I have solved this problem, but I doubt if I did this in the correct manner or even if the problem is correct.
Here is the problem as stated: [(37*8.9)(4-2(3-1))]-5(2/(2-7))(5)= ?
This is my work below:
[(37*8.9)(4-2(3-1))]-5(2/(2-7))(5)= ? (original problem)
1.) 37*8.9 = 329.3 (I start by going left to right - multiplying the first set of parentheses)
2.) (3-1) = 2 (I than remember to always work the inner most part of the problem)
3.) (4-2+2) = 2 (According to my instructor, when it comes to multiplication/division and addition/subtraction, one is able to go from left to right rather than adding or subtracting first. I am not sure if this is correct, but here this is done by his teaching)
4.) (2-7)=-5 (I than move on to the right side of the equal sign)
5.) 2/-5=-0.4 (I than finish what is inside of the parentheses)
6.) 329.3+4-5(-0.4)(5) (This is what I have thus far into the equation)
7.) -5(-0.4)=2
329.3+4+2(5)
8.) 2(5)=10
329.3+4+10
9.) 343.3 (After I combine the last 3 terms)
So... I'm pretty sure that there is something I missed while working the problem. Please help me understand what I might have done wrong. Thanks!




Answer by richwmiller(17219)

I got 10 as my answer which you have to admit is a nice small even number.
I would not do it left to right.
Follow pemdas rules of order of operations
I would also hold off on making numbers bigger. There is often a way to cancel things if you leave them in factored form
Actually let's do the right hand side
5(2/(2-7))(5)
2-7=-5
-5*2/-5*5
=10
(4-2(3-1))
3-1=2
4-2*2=0
I was right to hold off on multiplying
anything times 0 =0 so we are left with ten
so how you changed 4-2*2 in4-2+2
(4-2+2) = 2 so instead of zero you got 2